OpenStudy (inhaleof):

Choose the equation of the line passing through the point (-3, 1) and parallel to y = -x - 4.

OpenStudy (inhaleof):

y = x + 2 y = -x - 2 y = -x + 4 y = x - 4 -the options

OpenStudy (jonnyvonny):

For a function to be parallel to another, the slopes MUST be the same, and y-intercept must be different; in this case, the slope is -x and the y-intercept is -4. Since the slope has to be equal to the original function, and we're given the point (-3,1) we may:\[y=-x+b \rightarrow 1=-(-3)+b \rightarrow 1=3+b\]\[-2=b\] Plug that in now into y=mx+b. y=-x-2 b

OpenStudy (inhaleof):

Thanks alot (: